### 06 Mar 2026 v1

# Enhancing Neural Video Compression of Static Scenes with Positive-Incentive Noise

Abstract
Static scene videos, such as surveillance feeds and videotelephony streams, constitute a dominant share of storage consumption and network traffic. However, both traditional standardized codecs and neural video compression (NVC) methods struggle to encode these videos efficiently due to inadequate usage of temporal redundancy and severe distribution gaps between training and test data, respectively. While recent generative compression methods improve perceptual quality, they introduce hallucinated details that are unacceptable in authenticity-critical applications. To overcome these limitations, we propose to incorporate positive-incentive noise into NVC for static scene videos, where short-term temporal changes are reinterpreted as positive-incentive noise to facilitate model finetuning. By disentangling transient variations from the persistent background, structured prior information is internalized in the compression model. During inference, the invariant component requires minimal signaling, thus reducing data transmission while maintaining pixel-level fidelity. Preliminary experiments demonstrate a 73% Bjøntegaard delta (BD) rate saving compared to general NVC models. Our method provides an effective solution to trade computation for bandwidth, enabling robust video transmission under adverse network conditions and economic long-term retention of surveillance footage.
1 Introduction
With the expanding adoption of visual sensing and real-time communication, static scene videos, exemplified by surveillance feeds and videotelephony streams, constitute a dominant portion of storage consumption and network traffic. These videos typically contain limited and localized temporal variations with nearly stationary backgrounds, yet their prolonged duration results in immense data volumes, posing significant challenges to existing compression techniques. Traditional standardized video codecs, such as H.264/AVC (H264), H.265/HEVC (H265), and H.266/VVC (H266), exhibit insufficient coding efficiency for static scene videos, as their hand-crafted motion compensation and transform coding pipelines fail to fully exploit the temporal redundancy and structural regularity inherent in these videos.
To address this limitation, neural video compression (NVC) has achieved superior rate-distortion performance by leveraging end-to-end optimization and learned entropy models (ssf2020; liu2024bidirectional; DCVC-FM; DCVC-RT). However, their efficiency degrades drastically when deployed on static scene videos, primarily due to the severe distribution gap between the dynamic-rich training data and the static-dominant testing videos. This discrepancy leads to suboptimal motion modeling and inefficient bit allocation. Inspired by recent advances in video generation models, generative methods are employed to synthesize high-frequency details, thus eliminating compression artifacts and enhancing perceptual quality (GVC_teleai_2507; GVC_teleai_2512; GLC-video). Unfortunately, these approaches often introduce artificial textures and hallucinated details that are intolerable in critical scenarios such as surveillance and video conferencing, where pixel-level fidelity and content authenticity are mandatory requirements.
Beyond video compression, positive-incentive noise has demonstrated remarkable effectiveness across multiple machine learning tasks (PiNoise_li2022; PiNoise_zhang2025). By deliberately injecting structured perturbations during representation learning, positive-incentive noise encourages models to learn more discriminative features, thus improving robustness and generalization capability. The efficacy of this paradigm has been validated in diverse domains, including but not limited to visual-language alignment (PiNoise_VL_zhang2025; PiNoise_VLM_zhu2025), graph learning (PiNoise_graph_zhang2025; PiNoise_graph_contrastive_zhang2025), and unsupervised clustering (PiNoise_cluster_chen2024).
Motivated by these successes, we introduce positive-incentive noise into NVC methods for static scenes, where short-term temporal changes, such as object movements or light flickerings, are treated as positive-incentive noise during training. These dynamic elements incentivize the model to disentangle transient variations from persistent background representations, from which the structured prior information is extracted. With the assistance of positive-incentive noise, the prior knowledge can be progressively learned and implicitly stored in the compression model through on-device finetuning. During inference, the learned background prior requires minimal signaling, and the dynamic variations account for the primary components in the encoded and transmitted symbols. Our framework leverages positive-incentive noise as an instructive training mechanism instead of a generative solution, thus enhancing rate-distortion (RD) performance without compromising reconstruction authenticity.
The proposed framework aligns with the rapid development of artificial intelligence (AI) services deployed in communication and computation integrated networks, exemplified by the three-tier architecture described in the AI Flow framework (aiflow; aiflow_perspectives). Contrary to the mainstream cloud-centric computing scheme, these systems coordinate device, edge, and cloud resources to support latency-sensitive applications even in the most challenging circumstances, as explored by task-oriented communication for computer vision (shao2021learning; shao2024task; TOFC) and the law of information capacity in large language models (yuan2026informationcapacityevaluatingefficiency). The soaring capabilities of edge hardware have opened up the possibility of trading computation for bandwidth, where neural networks are employed on user devices to extract compact representations and minimize data transmission (fan2025computationbandwidthmemorytradeoffsunifiedparadigm). This trade-off exhibits particular significance when channel conditions fluctuate unpredictably or when spectrum resources become stringent due to high concurrency. Our method extends this idea to the scenario of static scene video compression and provides a robust solution to deliver smooth and high-definition video streams in adverse network environments. Apart from real-time communication, this approach also substantially reduces the storage costs of surveillance footage, rendering data retention over an extended period more economically viable.
2 Results
2.1 Implementation Details
We have recorded a total of over 132 hours of static scene surveillance videos in three different scenarios with a resolution of $2560\times 1440$ and a frame rate of 25. We finetune the space-scale flow (SSF) video compression model (ssf2020) on a 60-hour subset of this dataset using the codebase and checkpoints released by the CompressAI package (compressai). Five checkpoints corresponding to quality levels of 1, 4, 6, 7, 8 are utilized as initial weights, since they constitute the upper envelope of the RD performance curve before finetuning. During training, a 20-frame clip is randomly selected from each video sample with a resolution of $2560\times 1408$ to accommodate the network requirements, and the initial learning rate is set to $2\times 10^{-5}$. The remaining hyperparameters and implementations are identical to those of the CompressAI library, including the Adam optimizer, the reduce-on-plateau learning rate scheduling, an individual optimizer with a constant learning rate of $10^{-3}$ for the quantiles in the learned entropy models, the value of the Lagrange multiplier that controls the RD trade-off for each quality level, and using the mean squared error (MSE) loss to evaluation distortion.
2.2 RD Performance
We evaluate the SSF models on 100 randomly selected 600-frame clips outside the subset used for finetuning. The data size is measured by bits per pixel (BPP), and the reconstruction quality is assessed by peak signal-to-noise ratio (PSNR) in the YUV color space. The PSNR value of the Y channel is given a weighting coefficient six times larger than that of other channels, following the official implementation of deep contextual video coding (DCVC) evaluations. As illustrated in Figure 1, the proposed approach simultaneously increases the PSNR quality and reduces the data size for all five checkpoints, demonstrating the efficacy of positive-incentive noise in training video compression models. To achieve a PSNR quality lower than 42.5 dB, the proposed method can decrease the required BPP by at least 0.1, more than half of the original data volume. In a higher PSNR range, the rate savings are even more prominent, and the highest achievable PSNR is further increased from 44.92 dB to 48.84 dB while consuming only 37.6% of the original storage space. Overall, a 73.0% Bjøntegaard delta (BD) rate saving is achieved in comparison to original checkpoints, measured with cubic spline interpolation across a wide PSNR range from 32 dB to 45 dB.
Figure 1: RD performance comparison of the SSF model before and after finetuning. The proposed approach achieves a significant BD-rate reduction of 73%, demonstrating the efficacy of positive-incentive noise in training video compression models.
2.3 Visual Comparison
Figure 2 provides a visual comparison of video frames reconstructed by the traditional codec H264 and SSF models before and after finetuning. To ensure a fair comparison, we carefully select the hyperparameters of H264 and the checkpoints of SSF models to maintain a consistent data rate of approximately 0.2 BPP. The proposed approach based on positive-incentive noise enhances the PSNR quality from 38.70 dB to 46.27 dB with a slight reduction in BPP from 0.205 to 0.197. Direct visual comparison verifies the improved reconstruction quality, where the baseline model produces noticeable colored fringes on the text of the fire extinguisher container at this bitrate, while our finetuned model successfully preserves the original clarity. The actual reconstruction quality of the SSF model evaluated on our static scene videos is substantially lower than that of the traditional H264 codec. This result is contrary to the original paper, which shows consistent performance gains over H264 when BPP exceeds 0.1, mainly due to the distribution gap between the training dataset and static scene videos used in our evaluation. Guided by positive-incentive noise, our online finetuning method successfully bridges this gap and achieves superior performance compared to H264.
Figure 2: Visual comparison of video frames reconstructed by H264 and SSF models before and after finetuning at a BPP of 0.2. The proposed approach based on positive-incentive noise enhances the PSNR quality from 38.70 dB to 46.27 dB under nearly identical data rates.
3 Conclusion
This report proposes to utilize positive-incentive noise in enhancing the efficiency of NVC for static scene videos, where transient temporal changes serve as positive-incentive noise that benefits model finetuning. These variations stimulate the model to learn robust prior information from the persistent background, which is internalized in model parameters and requires minimal data transmission during inference. This method represents a new way of trading computation for bandwidth in communication and computation integrated networks, filling the vacancy in authenticity-critical scenarios left by previous generative video compression methods. Preliminary experiments demonstrate that the proposed approach achieves a significant BD-rate reduction of 73%, which enables economic long-term retention of surveillance footage and robust video streaming against network fluctuations. In the near future, we will extend the experiments to state-of-the-art NVC methods and provide concrete theoretical foundations on the role of positive-incentive noise.